OpenStudy (anonymous):

72a^2-50b^2 (Factor ?)

OpenStudy (anonymous):

72a^2 - 50b^2 = 2*(36a^2 - 25b^2). This is a difference of two squares (36, 25, with roots 6, 5). = 2*(6a-5b)(6a+5b).

OpenStudy (anonymous):

I actually didn't see that at first for some reason. So I threw it into wolfram alpha and got it: http://www.wolframalpha.com/input/?i=72a^2-50b^2 It can be useful to see if there might be any chance of turning it into a nicer form.